# Multi-source-free Domain Adaptation via Uncertainty-aware Adaptive Distillation

### v1

@@ Abstract
Source-free domain adaptation (SFDA) alleviates the domain discrepancy among data obtained from domains without accessing the data for the awareness of data privacy. However, existing conventional SFDA methods face inherent limitations in medical contexts, where medical data are typically collected from multiple institutions using various equipment. To address this problem, we propose a simple yet effective method, named Uncertainty-aware Adaptive Distillation (UAD) for the multi-source-free unsupervised domain adaptation (MSFDA) setting. UAD aims to perform well-calibrated knowledge distillation from (i) model level to deliver coordinated and reliable base model initialisation and (ii) instance level via model adaptation guided by high-quality pseudo-labels, thereby obtaining a high-performance target domain model. To verify its general applicability, we evaluate UAD on two image-based diagnosis benchmarks among two multi-centre datasets, where our method shows a significant performance gain compared with existing works. The code will be available soon.
@@ 1 Introduction
Unsupervised domain adaptation (UDA) is a promising streamline of works to compensate for the distributional discrepancy[1]. It seeks to utilise existing transferable knowledge from labelled data drawn from one or more source domains to recognise unlabelled data in the target domain[2]. UDA has shown great success in a broad spectrum of downstream applications, including classification [3] [4] [5], segmentation [6] [7] [8] and object detection [9] [10] by mitigating this domain shift.
Despite its great promises in general visual perception tasks, existing UDA approaches inherently fall short in medical scenarios where additional regulations on data sharing restrictions. To address the problems on medical images, source-free DA methods [4] have been developed, providing the pre-trained source model only instead of directly accessing the source data to preserve the privacy issue.
In this work, we investigate multi-source-free unsupervised domain adaptation (MSFDA) [11] [12] and improve the typical SFDA settings [4] [13] by introducing multiple source domains. It therefore holds the potential to serve as an appealing solution for real-world large-scale medical image analysis studies involving multiple centres. Several recent efforts have been made [12] [14] with preliminary attempts to the self-supervised clustering pseudo-labelling method [15], which is commonly adopted for MSFDA. However, they tend to be suboptimal particularly for medical image processing. Since the distinctions of the data from multiple centres are large, the models trained on datasets derived from single or multiple healthcare institutions have not demonstrated a consistent ability to generalise their applicability to external sites [16].
@@ 2.2 Uncertainty-aware Adaptive Distillation
In the proposed framework, we transfer the knowledge from multiple source models to adapt the target domain with pseudo-labels generated by distilling the proper source model. Technically, we learn a set of uncertainty (or its opposite, confidence) measures for both overall domain-wise and individual instance-wise distillation corresponding to each source model in the source model zoo. It evaluates the distributional distance of certain source models working on the target domain dataset and the quality of pseudo-labelling. Specifically, we introduce margin, defined as the difference between the predicted probabilities of the first and second most probable classes [17], as the metric to estimate the confidence measure:
$\mathcal{M}=\operatorname*{Topk}_{k=1}(\delta(\theta(x)))-\operatorname*{Topk}_{k=2}(\delta(\theta(x))),$ (1)
where $\delta(\cdot)$ denotes the Softmax Layer operation with $\delta_{j}(v)=\frac{\exp(v_{j})}{\sum_{i=1}^{K}\exp(v_{i})}$ for $j=1,...,K$ and $v\in\mathbb{R}^{K}\mapsto(0,1)^{K}$. Intuitively, if a model $\theta$ has a larger value of the margin $\mathcal{M}$ while predicting an instance, it is regarded as more optimal to extract the instance’s feature and finally does the classification task.
@@
Model-level UAD: In previous work related to multi-source domain adaptation[11], it was a common practice to involve all source models with varying weights in the subsequent fine-tuning stage. However, we found that if there is a significant domain gap between a particular source model and the target domain, negative transfer [18] could be incurred which results in biased adaptation. Thus, to initialise a base target model with minimal disturbance, we collect all pre-trained source models from each domain and estimate the overall confidence measure of each source model for predicting the target domain data. Specifically, for assessing the confidence of a source model $\theta_{S}^{j}$’s inference results on the target domain data, we average all confidence measures estimated for each instance of the target domain data as follows: $\mathcal{M}_{\text{j}}=\frac{\sum_{i=1}^{n_{T}}\mathcal{M}_{\text{i}}}{n_{T}}$. The source model with the largest confidence measure which is defined as $\varepsilon$ for the target domain, $\theta_{S}^{\ast}$, is regarded as the model conforming to the underlying data distribution closest to the target domain and can be considered as the optimal teacher:
$\varepsilon=\text{arg}\ {\text{max}}([\mathcal{M}_{\text{j}}]_{j=1}^{N}).$ (2)
@@
$\varepsilon_{i}=\text{arg}\ {\text{max}}([\mathcal{M}_{\text{i}}]_{i=1}^{n_{T}}),$ (3)
@@ 2.3 Temperature Scaling
In certain models, domain shift and limited data in source domains may result in over- and under-confidence in predicting target domain data which potentially triggers a mismatch between model prediction accuracy and confidence [19]. In other words, when this phenomenon occurs, the confidence measure $\varepsilon$ will no longer be an optimal measure for improving model prediction accuracy.
To address this problem, we embedded Temperature Scaling (TS) which acts on prediction probabilities to calibrate the logits prior to confidence measurement. In our approach, TS is capable of effectively regularising the representation of uncertainty in model predictions, and a more precise and unbiased representation of uncertainty is preferable for the process of knowledge distillation. The parameter $\mathcal{T}$ is the so-called temperature, which yields softer probability estimates with larger a temperature to alleviate over-confidence in the model. For every source model $[\theta_{S}^{j}]_{j=1}^{N}$, we learn $\mathcal{T}_{\text{j}}$ by setting an initialisation value $\mathcal{T}_{\text{initial}}$ and applying temperature scaling on the target domain data $\mathcal{D}_{T}$: $\mathcal{T}_{\text{j}}=\text{TS-Alg}([\theta_{S}^{j}]_{j=1}^{N},\mathcal{D}_{T})$. Specifically, the temperature scaling models are tuned by minimising expected calibration error (ECE), a.k.a., calibration gap, which is defined as the difference between accuracy and confidence for a given bin [20]:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n_{T}}\Big{|}\text{acc}(B_{m})-\text{conf}(B_{m})\Big{|},$ (6)
@@ 3.1 Dataset and Implementation Details
  • Multi-centre Diabetic Retinopathy (DR) dataset: The multi-centre DR dataset, which measures DR grades (no DR, mild DR, moderate DR, severe DR and proliferative DR), consists of three public datasets (domains) APTOS 2019 [21], DDR [22], and IDRiD [23] with counts $3660$, $13673$, and $516$ correspondingly.
  • Skin Cancer MNIST HAM10000 [24]: To investigate the classification of lesions as benign or malignant in different parts of the human body, we split it into four domains by skin lesion locations which are back, face, lower extremity, and upper extremity with counts $2192$, $745$, $2077$ and $1118$ respectively.
@@
AaD (22’) [13] 36.13 33.07 46.32 38.51 64.55 64.30 65.14 72.36 66.59
DECISION (21’) [11] 57.32 45.43 58.33 53.69 74.27 76.24 71.06 78.98 75.14
CAiDA (21’) [12] 71.74 44.98 50.97 55.90 73.68 73.83 79.59 78.80 76.48
@@
Implementation Details: Following the top-rank solution for medical image classification [25], we employ DenseNet-121 as the backbone. In the source model training process, we use smooth labels instead of the usual one-hot labels to reduce overfitting and label noise. The maximum number of epochs $\mathcal{N}_{\text{epoch}}$ for both DR and HAM10000 datasets is set to $100$; while during the UAD process, the $\mathcal{N}_{\text{epoch}}$ is set to $15$ with a series of updated pseudo-labels at the start of each. The batch size is set to 32. For each epoch, there are $\mathcal{N}_{\text{training data}}/32$ iterations in domains. We use $\mathcal{T}_{\text{initial}}=\log{(1/1.5)}$ and $1.5$ for the DR dataset and the HAM10000 dataset, respectively. For both source models pre-training and adaptive distillation, we leverage stochastic gradient descent with momentum value $0.9$ and weight decay $10^{-3}$, with the learning rate scheduling method [3] during the model learning progress.
@@ 3.2 Comparison Experiments
For experimental comparison, we included one existing SFDA framework AaD [13] with multi-source extension and two MSFDA frameworks DECISION [11] and CAiDA [12] as baseline methods. We re-implement them following their default settings. The experimental results are reported in Table 1. The multi-source extension of AaD is implemented via an ensemble that passes the target data through each of the adapted source model and takes an average of the soft prediction to obtain the test label. By exploring the experimental results of iterations during the SFDA process for DECISION, we noted that, except for the target domain I in DR and F in HAM10000, the performance of the DECISION model deteriorates as the iterations increase for training the target model. This phenomenon is also observed in the CAiDA framework, although the degradation in model performance in the domain adaptation process is not as severe as in the DECISION framework. Intuitively, in a domain-biased and unsupervised setting, the model overfits to noisy labels when training on the target data. It is due to the effect of the involvement of inappropriate source models and low-quality pseudo-labels generated.
@@ 5 Compliance with Ethical Standards
This research study was conducted retrospectively using human subject data made available in open access by [21, 22, 23, 24]. Ethical approval was not required as confirmed by the license attached with the open access data.

### v2

@@ Abstract
Source-free domain adaptation (SFDA) alleviates the domain discrepancy among data obtained from domains without accessing the data for the awareness of data privacy. However, existing conventional SFDA methods face inherent limitations in medical contexts, where medical data are typically collected from multiple institutions using various equipment. To address this problem, we propose a simple yet effective method, named Uncertainty-aware Adaptive Distillation (UAD) for the multi-source-free unsupervised domain adaptation (MSFDA) setting. UAD aims to perform well-calibrated knowledge distillation from (i) model level to deliver coordinated and reliable base model initialisation and (ii) instance level via model adaptation guided by high-quality pseudo-labels, thereby obtaining a high-performance target domain model. To verify its general applicability, we evaluate UAD on two image-based diagnosis benchmarks among two multi-centre datasets, where our method shows a significant performance gain compared with existing works. The code is available at https://github.com/YXSong000/UAD.
@@ 1 Introduction
Unsupervised domain adaptation (UDA) is a promising streamline of works to compensate for the distributional discrepancy[Fan_2023_ICCV]. It seeks to utilise existing transferable knowledge from labelled data drawn from one or more source domains to recognise unlabelled data in the target domain[ben-david_theory_2010]. UDA has shown great success in a broad spectrum of downstream applications, including classification [pmlr-v37-ganin15] [liang2020we] [li2022domain], segmentation [liu2020unsupervised] [wang2022cris] [fan2024learning] and object detection [hsu2020progressive] [Liu2022DecomposeTA] by mitigating this domain shift.
Despite its great promises in general visual perception tasks, existing UDA approaches inherently fall short in medical scenarios where additional regulations on data sharing restrictions. To address the problems on medical images, source-free DA methods [liang2020we] have been developed, providing the pre-trained source model only instead of directly accessing the source data to preserve the privacy issue.
In this work, we investigate multi-source-free unsupervised domain adaptation (MSFDA) [Ahmed_2021_CVPR] [NEURIPS2021_Dong] and improve the typical SFDA settings [liang2020we] [yang2022attracting] by introducing multiple source domains. It therefore holds the potential to serve as an appealing solution for real-world large-scale medical image analysis studies involving multiple centres. Several recent efforts have been made [NEURIPS2021_Dong] [han2023discriminability] with preliminary attempts to the self-supervised clustering pseudo-labelling method [wang2023mosaic], which is commonly adopted for MSFDA. However, they tend to be suboptimal particularly for medical image processing. Since the distinctions of the data from multiple centres are large, the models trained on datasets derived from single or multiple healthcare institutions have not demonstrated a consistent ability to generalise their applicability to external sites [zech2018variable].
@@ 2.2 Uncertainty-aware Adaptive Distillation
In the proposed framework, we transfer the knowledge from multiple source models to adapt the target domain with pseudo-labels generated by distilling the proper source model. Technically, we learn a set of uncertainty (or its opposite, confidence) measures for both overall domain-wise and individual instance-wise distillation corresponding to each source model in the source model zoo. It evaluates the distributional distance of certain source models working on the target domain dataset and the quality of pseudo-labelling. Specifically, we introduce margin, defined as the difference between the predicted probabilities of the first and second most probable classes [Settles2009ActiveLL], as the metric to estimate the confidence measure:
$\mathcal{M}=\operatorname*{Topk}_{k=1}(\delta(\theta(x)))-\operatorname*{Topk}_{k=2}(\delta(\theta(x))),\$(1)
where $\delta(\cdot)$ denotes the Softmax Layer operation with $\delta_{j}(v)=\frac{\exp(v_{j})}{\sum_{i=1}^{K}\exp(v_{i})}$ for j=1,,Kj=1,...,K and $v\in\mathbb{R}^{K}\mapsto(0,1)^{K}$. Intuitively, if a model $\theta$ has a larger value of the margin $\mathcal{M}$ while predicting an instance, it is regarded as more optimal to extract the instance’s feature and finally does the classification task.
@@
Model-level UAD: In previous work related to multi-source domain adaptation[Ahmed_2021_CVPR], it was a common practice to involve all source models with varying weights in the subsequent fine-tuning stage. However, we found that if there is a significant domain gap between a particular source model and the target domain, negative transfer [8953565] could be incurred which results in biased adaptation. Thus, to initialise a base target model with minimal disturbance, we collect all pre-trained source models from each domain and estimate the overall confidence measure of each source model for predicting the target domain data. Specifically, for assessing the confidence of a source model $\theta_{S}^{j}$’s inference results on the target domain data, we average all confidence measures estimated for each instance of the target domain data as follows: $\mathcal{M}_{\text{j}}=\frac{\sum_{i=1}^{n_{T}}\mathcal{M}_{\text{i}}}{n_{T}}$. The source model with the largest confidence measure which is defined as $\varepsilon$ for the target domain, $\theta_{S}^{\ast}$, is regarded as the model conforming to the underlying data distribution closest to the target domain and can be considered as the optimal teacher:
$\varepsilon=\text{arg}\ {\text{max}}([\mathcal{M}_{\text{j}}]_{j=1}^{N}).\$(2)
@@
$\varepsilon_{i}=\text{arg}\ {\text{max}}([\mathcal{M}_{\text{i}}]_{i=1}^{n_{T}}),\$(3)
@@ 2.3 Temperature Scaling
In certain models, domain shift and limited data in source domains may result in over- and under-confidence in predicting target domain data which potentially triggers a mismatch between model prediction accuracy and confidence [NEURIPS2021_f8905bd3]. In other words, when this phenomenon occurs, the confidence measure $\varepsilon$ will no longer be an optimal measure for improving model prediction accuracy.
To address this problem, we embedded Temperature Scaling (TS) which acts on prediction probabilities to calibrate the logits prior to confidence measurement. In our approach, TS is capable of effectively regularising the representation of uncertainty in model predictions, and a more precise and unbiased representation of uncertainty is preferable for the process of knowledge distillation. The parameter $\mathcal{T}$ is the so-called temperature, which yields softer probability estimates with larger a temperature to alleviate over-confidence in the model. For every source model $\theta_{S}^{j}]_{j=1}^{N}$, we learn $\mathcal{T}_{\text{j}}$ by setting an initialisation value $\mathcal{T}_{\text{initial}}$ and applying temperature scaling on the target domain data $\mathcal{D}_{T}$: $\mathcal{T}_{\text{j}}=\text{TS-Alg}([\theta_{S}^{j}]_{j=1}^{N},\mathcal{D}_{T})$. Specifically, the temperature scaling models are tuned by minimising expected calibration error (ECE), a.k.a., calibration gap, which is defined as the difference between accuracy and confidence for a given bin [pmlr-v70-guo17a]:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n_{T}}\Big|\text{acc}(B_{m})-\text{conf}(B_{m})\Big|,\$(6)
@@ 3.1 Dataset and Implementation Details
  • Multi-centre Diabetic Retinopathy (DR) dataset: The multi-centre DR dataset, which measures DR grades (no DR, mild DR, moderate DR, severe DR and proliferative DR), consists of three public datasets (domains) APTOS 2019 [aptos2019-blindness-detection], DDR [li2019diagnostic], and IDRiD [data3030025] with counts $3660$, $13673$, and $516$ correspondingly.
  • Skin Cancer MNIST HAM10000 [tschandl2018ham10000]: To investigate the classification of lesions as benign or malignant in different parts of the human body, we split it into four domains by skin lesion locations which are back, face, lower extremity, and upper extremity with counts $2192$, $745$, $2077$ and $1118$ respectively.
@@
AaD (22’) [yang2022attracting] 36.13 33.07 46.32 38.51 64.55 64.30 65.14 72.36 66.59
DECISION (21’) [Ahmed_2021_CVPR] 57.32 45.43 58.33 53.69 74.27 76.24 71.06 78.98 75.14
CAiDA (21’) [NEURIPS2021_Dong] 71.74 44.98 50.97 55.90 73.68 73.83 79.59 78.80 76.48
@@
Implementation Details: Following the top-rank solution for medical image classification [liu2021federated], we employ DenseNet-121 as the backbone. In the source model training process, we use smooth labels instead of the usual one-hot labels to reduce overfitting and label noise. The maximum number of epochs $\mathcal{N}_{\text{epoch}}$ for both DR and HAM10000 datasets is set to $100$; while during the UAD process, the $\mathcal{N}_{\text{epoch}}$ is set to $15$ with a series of updated pseudo-labels at the start of each. The batch size is set to 32. For each epoch, there are $\mathcal{N}_{\text{training data}}/32$ iterations in domains. We use $\mathcal{T}_{\text{initial}}=\log{(1/1.5)}$ and $1.5$ for the DR dataset and the HAM10000 dataset, respectively. For both source models pre-training and adaptive distillation, we leverage stochastic gradient descent with momentum value $0.9$ and weight decay $10^{-3}$, with the learning rate scheduling method [pmlr-v37-ganin15] during the model learning progress.
@@ 3.2 Comparison Experiments
For experimental comparison, we included one existing SFDA framework AaD [yang2022attracting] with multi-source extension and two MSFDA frameworks DECISION [Ahmed_2021_CVPR] and CAiDA [NEURIPS2021_Dong] as baseline methods. We re-implement them following their default settings. The experimental results are reported in Table 1. The multi-source extension of AaD is implemented via an ensemble that passes the target data through each of the adapted source model and takes an average of the soft prediction to obtain the test label. By exploring the experimental results of iterations during the SFDA process for DECISION, we noted that, except for the target domain I in DR and F in HAM10000, the performance of the DECISION model deteriorates as the iterations increase for training the target model. This phenomenon is also observed in the CAiDA framework, although the degradation in model performance in the domain adaptation process is not as severe as in the DECISION framework. Intuitively, in a domain-biased and unsupervised setting, the model overfits to noisy labels when training on the target data. It is due to the effect of the involvement of inappropriate source models and low-quality pseudo-labels generated.
@@ 5 Compliance with Ethical Standards
This research study was conducted retrospectively using human subject data made available in open access by [aptos2019-blindness-detection, li2019diagnostic, data3030025, tschandl2018ham10000]. Ethical approval was not required as confirmed by the license attached with the open access data.